i really need your help on this i tried but ............ the question says share $873 between Peter and John so that Peter gets 3 1/2 times more how much does each get pliz im beggin
Found 2 solutions by Earlsdon, colliefan:
Answer by Earlsdon(6294)

Let P = the amount that Peter will get while J = the amount that John will get.
We know that these two amounts add up to $873.00, so you can write:
1) P+J = 873
Now it says that Peter gets 3.5 time more than John so this can be expressed as:
2) P = 3.5J Now substitute this into equation 1) to get:
1a) 3.5J+J = 873 Solve this for J.
1b) 4.5J = 873 Divide by 4.5
1c) J = 194
So John gets $194.00 and Peter gets 3.5($194) = $679.00

Answer by colliefan(242)

Call the amount that John gets x.
Then the amount that Peter gets is 3.5*x
The other information gives the equation:
x + 3.5*x = 873
(remember that the units are dollars)
4.5 x = 873
x=873/4.5
x = 194 = John's amt
Peter's amt is 3.5 times 194.